### v1

@@ 2.2 The Cooperate Games
It yields the exact form of the optimal control
$\check{v}_{t}^{i}=-\frac{2a(t)\check{X}_{t}^{i}+b(t)+\bar{p}(t)+\alpha l(t)}{c}.$ (2.23)
Applying (2.23), we simplify the HJB equation into the following ODE system
$\qquad\qquad\qquad a^{\prime}(t)+\frac{\eta}{2}-\frac{2a^{2}(t)}{c}=0,$ $a(T)=\frac{\gamma}{2},$ (2.24)
$\qquad\qquad\qquad b^{\prime}(t)-\eta\kappa-\frac{2a(t)(\bar{p}(t)+b(t)+\alpha l(t))}{c}=0,$ $b(T)=-\gamma\zeta,$ (2.25)
$\qquad\qquad\qquad f^{\prime}_{i}(t)+\sigma_{i}^{2}a(t)+\frac{\eta\kappa^{2}}{2}-\frac{(b(t)+\bar{p}(t)+\alpha l(t))^{2}}{2c}=0,$ $f_{i}(T)=\frac{\gamma\zeta^{2}}{2},$ (2.26)
$\qquad\qquad\qquad l^{\prime}(t)-\alpha l(t)-\bar{q}(t)=0,$ $l(T)=0.$ (2.27)
We state that (2.24),(2.25),(2.26) and (2.27) give a unique solution established on $[0,T]$. To prove this, we may argue as follows. First, (2.24) gives a unique solution on $[0,T]$, as a result of subsection 2.1. Then, similar as in the proof of subsection 2.1, (2.27) and (2.25) both give a unique solution on $[0,T]$. Finally, the existence and uniqueness of solutions to (2.26) is obvious, once we integrate both sides of (2.26) from $t$ to $T$. We will not verify the the optimality of the solution given by (2.23), for the optimality property is not necessary for us to arrive at the results in Theorem 2.11.
As in the discussion of the non-cooperate games section, if the agents adopt the strategies introduced in (2.23), the average of $X_{t}^{i}$, abbreviated as $\overline{X}_{t}$, can be approximated by a deterministic process $\bar{x}$, whose dynamics satisfy
@@
This implies that, the requirement in step 3 is equivalent to $\bar{q}(t)=-\frac{2a(t)\bar{x}(t)+b(t)+\bar{p}(t)+\alpha l(t)}{c}$. By integrating the process (2.21),(2.25), (2.27) and the above evolution law for $\bar{q}$, we observe that this is also equivalent to ensure that the following dynamics
@@
together with the corresponding initial or terminal conditions,
@@
holds true simultaneously.
Again, it is crucial that (2.28), together with (2.29), indeed gives a solution on $[0,T]$. Similar as in the previous case, the main difficulty in determining the existence of this solution is that the boundary values are specified at different points. By this observation, we first construct two processes whose initial values are only specified at $t=0$ and study their properties.
@@
with $(\bar{p}_{1}(0),\ \bar{x}_{1}(0),\ {b}_{1}(0),\ {l}_{1}(0))^{\top}=(0,0,1,0)^{\top}$ and $(\bar{p}_{2}(0),\ \bar{x}_{2}(0),\ {b}_{2}(0),\ {l}_{2}(0))^{\top}=(0,0,0,1)^{\top}$, give a unique solution, denoted by $(\phi_{1}^{*}(t))_{t\in[0,T]}=((\bar{p}_{1}(t),\ \bar{x}_{1}(t),\ b_{1}(t),\ l_{1}(t))^{\top})_{t\in[0,T]}$ and $(\phi_{2}^{*}(t))_{t\in[0,T]}$ $((\bar{p}_{2}(t),\ \bar{x}_{2}(t),\ b_{2}(t),\ l_{2}(t))^{\top})_{t\in[0,T]}$ respectively.
Similar as in the non-cooperate case, we will show that, the solution to (2.28), if given initial values specified only at $t=0$, must give a unique solution.
@@
subsection 2.2 states that, given $\bar{p}(0)=p_{0}$ and $\bar{x}(0)=\bar{x}_{0}^{N}$, the solution to (2.28) is uniquely determined by $b(0)$ and $l(0)$. Now, we propose a sufficient condition for the existence and uniqueness for (2.28) and (2.29), as shown in Assumption $\boldsymbol{(A_{3})}$.
Assumption $\boldsymbol{(A_{3}):}$ The processes $((\bar{p}_{1}(t),\ \bar{x}_{1}(t),\ b_{1}(t),\ l_{1}(t))^{\top})_{t\in[0,T]}$ and $((\bar{p}_{2}(t),\ \bar{x}_{2}(t)$, $b_{2}(t),$ $l_{2}(t))^{\top})_{t\in[0,T]}$, which are given by (2.30) and (2.31) respectively, satisfy that $({b}_{1}(T),\ {l}_{1}(T))^{\top}$, $({b}_{2}(T),\ {l}_{2}(T))^{\top}$ are linearly independent vectors.
@@
If Assumption $\boldsymbol{(A_{3})}$ holds true, then the existence and uniqueness of (2.28) and (2.29) hold true.
As in other MFG works, We need coercive conditions to prevent the absence of an optimal or $\epsilon$-optimal control. However, this condition may not hold for all positive numbers $c$ and $\eta$. This drives us to propose the following assumption to ensure that the coercive condition holds true.
@@
To see why the assumption is necessary, consider the following case. If $v_{t}^{i}$ remains being a large constant, then $\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\sum_{i=1}^{N}(v_{t}^{i}P_{t})dt\right]$ grows into a very small negative number that may exceed the influence of the positive terms $\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\sum_{i=1}^{N}\left(\frac{c}{2}{(v_{t}^{i})}^{2}+\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}\right)dt\right]$ when $c$ and $\eta$ are sufficiently small. Therefore, $J_{soc}$ can be arbitrarily small as the norm of the control increases, which makes the coercive condition fail, and the discussion of the problem meaningless. Note that the coefficient of the term $v_{t}^{i}P_{t}$ is 1 within $J_{soc}$, we posit that $\frac{c}{2}>1$ and proper lower bound for $\eta$ would suffice to guarantee coercivity; this will be rigorously shown in Lemma 2.5 below.
Given the aforementioned assumptions, we propose the process as shown in Theorem 2.11, and demonstrate that it provides a set of $\epsilon$-optimal strategies.
@@
Assume that Assumption $\boldsymbol{(A_{2})}$, $\boldsymbol{(A_{3})}$ and $\boldsymbol{(A_{4})}$ hold, and the unique solution for (2.28) and (2.29) is $(\bar{p}(t),\ \bar{x}(t),\ b(t),\ l(t))^{\top}$. We construct the process
@@
Then, we claim that the set of strategies $\{\check{v}^{i}\}_{1\leq i\leq N}$ is a set of $\epsilon$-optimal strategies for $J_{soc}$. Moreover, specific calculation shows $\epsilon_{N}=O(\frac{1}{\sqrt{N}})$ in (2.5).
Assume there exists an alternative set of admissible control strategies $\boldsymbol{v}=\{v^{i}\}_{1\leq i\leq N}$, which differs from $\check{\boldsymbol{v}}=\{\check{v}^{i}\}_{1\leq i\leq N}$. The corresponding processes it induces evolve as follows
@@
Our objective is to show that (2.5) holds true, or equivalently
@@
where the $\epsilon_{N}$ term is an $O(\frac{1}{\sqrt{N}})$ number. To do this, we introduce two lemmas. Lemma 2.5 is proposed to characterize the coercive condition. However, to facilitate a clear demonstration of the coercivity, we begin by presenting Lemma 2.4.
@@
We construct the process
@@
Now, we claim that, for any stochastic control $v\in\mathcal{A}$, the following result
@@
holds true.
We now proceed to the rigorous formulation of the coercivity condition, as shown in Lemma 2.5.
@@
Assume that Assumptions $\boldsymbol{(A_{2})}$ and $\boldsymbol{(A_{4})}$ hold true. Then there exists constants $\epsilon>0,\ \epsilon^{\prime}>0$, and $C\in\mathbb{R}$, all independent of $N$, such that
@@
Up to this point, we have done the preparations, so we proceed to the proof of Theorem 2.11.
@@
and $(\check{X}_{t}^{i},\ \check{P}_{t},\ \check{Q}_{t})$ are given by (2.32). Within these notions, the variation in $J_{soc}$ can be written as
@@
We transform the difference in terminal value into its integration form by using Ito’s formula:
@@
Plugging (2.35) into (2.34), we write
@@
We now provide a more detailed explanation of how we derive $I_{1}$ and what $I_{2}$ characterizes. To obtain $I_{1}$, we first approximate $\sum_{i=1}^{N}\check{v}_{t}^{i}\Delta P_{t}$ with $\sum_{i=1}^{N}\overline{q}(t)\Delta P_{t}$ and approximate $\sum_{i=1}^{N}\check{P}_{t}\Delta v_{t}^{i}$ with $\sum_{i=1}^{N}\overline{p}(t)\Delta v_{t}^{i}$ when calculating the integral of $\sum_{i=1}^{N}(\Delta P_{t}\check{v}_{t}^{i}+\check{P}_{t}\Delta v_{t}^{i})$. Then, we apply (2.35) to transform the terminal terms into the integral form. Following this, we use (2.32) to eliminate the $\check{v}_{t}^{i}$ term (i.e. turn it into the expression of $\check{X}_{t}^{i}$ and the continuous functions). Finally, we simply integrate the same terms to obtain $I_{1}$. Obviously, it characterizes how much approximately $J_{soc}$ increases after a deviation of the proposed process is observed. The second part, $I_{2}$, explicitly measures the error caused by this approximation.
Working exactly as in the proof of Lemma 2.1, we obtain that there exists an $O(1)$ constant, which we set as $K^{\prime\prime}$, that satisfies
@@
$|J_{i}(\check{v}^{i},\check{\boldsymbol{v}}^{-i})|\leq K^{\prime\prime},\quad i=1,2,\ldots,N.$
This leads to
@@
From the coercive condition given by Lemma 2.5, we conclude that there exists an $O(1)$ constant $M_{1}>0$ such that, if $\mathbb{E}\left[\int_{0}^{T}Q_{t}^{2}dt\right]>M_{1}$ is satisfied, then we have
@@
This inequality would give us the desired result. Therefore, it suffices to prove (2.33) under the case where $\mathbb{E}\left[\int_{0}^{T}Q_{t}^{2}dt\right]\leq M_{1}$. Indeed, under this case, we will demonstrate that
@@
First, we show that the first part holds. The next step relies on the following identity
@@
We will then demonstrate how it is proved. Ito’s formula, together with l{l}’s evolution process displayed in (2.27), implies that
@@
$=\mathbb{E}\left[\int_{0}^{T}\left(\bar{q}(t)\Delta u_{t}^{i}dt+\alpha{l}(t)\Delta v_{t}^{i}\right)dt\right].$ (2.38)
$\Delta{P_{t}}$ and $\Delta u_{t}^{i}$ evolve as follows,
@@
Note that $\Delta Q_{t}$ is the average of $\Delta v_{t}^{i}$, it is easily checked from (2.39) and (2.40) that
$\frac{1}{N}\sum_{i=1}^{N}\Delta u_{t}^{i}=-\Delta P_{t}.$ (2.41)
This gives
@@
where the first and the second equation are derived from (2.41) and (2.2) respectively. This proves (2.37). By a simple calculation, (2.37) simplifies (2.36) into
@@
Using Ito’s formula again, we obtain
@@
We introduce the following notation
$\overline{\Delta X_{t}}:=\frac{1}{N}\sum_{i=1}^{N}\Delta X_{t}^{i}.$
Note that
@@
$(\overline{\Delta X_{t}},\ \Delta P_{t})^{\top}$ follows the evolution process introduced in Lemma 2.4, so we conclude
@@
In order to use this inequality, we make use of Assumption $\boldsymbol{(A_{4})}$. Fundamental inequality implies
@@
As in the proof of Theorem 2.6, the convexity of $\frac{\eta}{2}x^{2}$ leads to
@@
where we apply Jensen’s inequality. Specifically, we substitute $\sum_{i=1}^{N}(\Delta X_{t}^{i})^{2}$ with the smaller (or equal) number $N\overline{\Delta X_{t}}^{2}$, and $\sum_{i=1}^{N}(\Delta v_{t}^{i})^{2}$ with $N\Delta Q_{t}^{2}$. Note that (2.42) gives a non-negative value, so
@@
The proof for the first part is completed, and it is straightforward to prove the second part. Now, similar as in Theorem 2.6’s proof, we have
@@
From the Cauchy-Schwartz inequality, the triangle inequality, and $\|Q\|^{2}\leq M_{1}$, we have
@@
We solve from (2.39),
$\Delta P_{t}=-\alpha\int_{0}^{t}e^{-\alpha(t-s)}\Delta Q_{s}ds.$
Cauchy-Schwartz theorem implies that
@@
We deduce from (2.43) and (2.44) that

### v2

@@ 4 The Cooperate Games
From this, we obtain the explicit optimal control
$\check{v}_{t}^{i}=-\frac{2a_{i}(t)\check{X}_{t}^{i}+b_{i}(t)+\bar{p}(t)+\alpha l_{i}(t)}{c}.$ (4.6)
Substituting (4.6) into the HJB equation yields the following ODE system:
$\qquad\qquad\qquad a^{\prime}_{i}(t)+\frac{\eta}{2}-\frac{2a^{2}_{i}(t)}{c}=0,$ $a_{i}(T)=\frac{\gamma}{2},$ (4.7)
$\qquad\qquad\qquad b^{\prime}_{i}(t)-\eta\kappa-\frac{2a_{i}(t)(\bar{p}(t)+b_{i}(t)+\alpha l_{i}(t))}{c}=0,$ $b_{i}(T)=-\gamma\zeta,$ (4.8)
$\qquad\qquad\qquad f^{\prime}_{i}(t)+\sigma_{i}^{2}a_{i}(t)+\frac{\eta\kappa^{2}}{2}-\frac{(b_{i}(t)+\bar{p}(t)+\alpha l_{i}(t))^{2}}{2c}=0,$ $f_{i}(T)=\frac{\gamma\zeta^{2}}{2},$ (4.9)
$\qquad\qquad\qquad l^{\prime}_{i}(t)-\alpha l_{i}(t)-\bar{q}(t)=0,$ $l_{i}(T)=0.$ (4.10)
We assert that (4.7)–(4.10) admit a unique solution on $[0,T]$. The proof is nearly identical to that of section 3, so we only provide a brief outline. First, (4.7) has a unique solution on $[0,T]$ by section 3. Next, as in the proof of section 3, both (4.10) and (4.8) admit unique solutions on $[0,T]$. Finally, uniqueness of the solution to (4.9) follows immediately by integrating the equation from $t$ to $T$. Since $a_{i}(t),\ b_{i}(t)$ and $l_{i}(t)$ are identical for all agents, we abbreviate them as a(t),b(t)a(t),\ b(t) and $l(t)$, respectively. Following the same argument as in Theorem 3.2, one can verify the optimality of (4.6), although such verification is not required for establishing the results in Theorem 4.5.
As in the analysis of the non-cooperative game, when all agents adopt the strategy in (4.6), the empirical mean of $X_{t}^{i}$, denoted by $\overline{X}_{t}$, can be approximated by a function $(\bar{x}(t))_{t\in[0,T]}$, whose dynamics satisfy
@@
This implies that the requirement in step 3 is equivalent to $\bar{q}(t)=-\frac{2a(t)\bar{x}(t)+b(t)+\bar{p}(t)+\alpha l(t)}{c}$. Combining this with (4.3), (4.8), (4.10), we observe that the requirement is equivalent to ensuring that the system
@@
together with the associated initial and terminal conditions,
@@
admits a solution on $[0,T]$.
As before, the main challenge lies in the boundary conditions being imposed at different time points. To overcome this, we first construct two processes whose initial values are both specified at $t=0$, as described below.
@@
associated with initial conditions $(\bar{p}_{1}(0),\ \bar{x}_{1}(0),\ {b}_{1}(0),\ {l}_{1}(0))^{\top}=(0,0,1,0)^{\top}$ and $(\bar{p}_{2}(0),\ \bar{x}_{2}(0),\ {b}_{2}(0),\ {l}_{2}(0))^{\top}=(0,0,0,1)^{\top}$, admit a unique solution, denoted by $(\phi_{1}^{*}(t))_{t\in[0,T]}=((\bar{p}_{1}(t),\ \bar{x}_{1}(t),\ b_{1}(t),\ l_{1}(t))^{\top})_{t\in[0,T]}$ and $(\phi_{2}^{*}(t))_{t\in[0,T]}$ $((\bar{p}_{2}(t),\ \bar{x}_{2}(t),\ b_{2}(t),\ l_{2}(t))^{\top})_{t\in[0,T]}$ respectively.
Next, we show that when the initial values of (4.11) are specified only at $t=0$, the system admits a unique solution. This implies that the solution to (4.11) is uniquely determined by $b(0)$ and $l(0)$, once given $\bar{p}(0)=p_{0}$ and $\bar{x}(0)=\bar{x}_{0}^{N}$.
@@
We now introduce a sufficient condition ensuring the existence and uniqueness of solutions to (4.11)–(4.12), stated in Assumption $\boldsymbol{(A_{3})}$.
Assumption $\boldsymbol{(A_{3}):}$ The systems $((\bar{p}_{1}(t),\ \bar{x}_{1}(t),\ b_{1}(t),\ l_{1}(t))^{\top})_{t\in[0,T]}$ and $((\bar{p}_{2}(t),\ \bar{x}_{2}(t)$, $b_{2}(t),$ $l_{2}(t))^{\top})_{t\in[0,T]}$, which are given by (4.13) and (4.14) respectively, satisfy that $({b}_{1}(T),\ {l}_{1}(T))^{\top}$, $({b}_{2}(T),\ {l}_{2}(T))^{\top}$ are linearly independent vectors.
section 4 below illustrates that Assumption $\boldsymbol{(A_{3})}$ is sufficient for ensuring the existence of solution to (4.11) and (4.12).
@@
If Assumption $\boldsymbol{(A_{3})}$ holds, then the existence and uniqueness of solution to (4.11) and (4.12) also hold.
As in other works on mean field games, a coercivity condition is required to prevent the absence of an optimal or $\epsilon$-optimal control. However, this condition need not hold for arbitrary positive parameters $c$ and $\eta$. To ensure coercivity, we impose the following assumption.
@@
To illustrate the necessity of this assumption, consider the case in which $v_{t}^{i}$ remains being a large constant. Then, the term $\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\sum_{i=1}^{N}(v_{t}^{i}P_{t})dt\right]$ can become a large negative number, potentially dominating the positive contributions from $\frac{1}{N}\mathbb{E}\left[\int_{0}^{T}\sum_{i=1}^{N}\left(\frac{c}{2}{(v_{t}^{i})}^{2}+\frac{\eta}{2}(X_{t}^{i}-\kappa)^{2}\right)dt\right]$ when $c$ and $\eta$ are sufficiently small. Consequently, $J_{soc}$ may approach -$\infty$ as the control norm increases, violating coercivity. Since the coefficient in front of $v_{t}^{i}P_{t}$ in $J_{soc}$ is 1, requiring $\frac{c}{2}>1$ and a proper lower bound for $\eta$ is sufficient to guarantee coercivity, as stated in Assumption $\boldsymbol{(A_{4})}$.
Under these assumptions, we construct the process described in Theorem 4.5 and show that it yields a set of $\epsilon$-optimal strategies.
@@
Assume that Assumption $\boldsymbol{(A_{2})}$, $\boldsymbol{(A_{3})}$ and $\boldsymbol{(A_{4})}$ hold, and the unique solution for (4.11) and (4.12) is $(\bar{p}(t),\ \bar{x}(t),\ b(t),\ l(t))^{\top}$. We construct the system
@@
Then, the set of strategies $\{\check{v}^{i}\}_{1\leq i\leq N}$ is a set of $\epsilon$-optimal strategies for $J_{soc}$. Moreover, specific calculation shows that under $\{\check{v}_{t}^{i}\}_{i=1}^{N}$, $\epsilon_{N}=O(\frac{1}{\sqrt{N}})$ in (2.5).
Suppose there exists another admissible control profile $\boldsymbol{v}=\{v^{i}\}_{1\leq i\leq N}$, and let the corresponding processes evolve according to
@@
Our objective is to show that (2.5) holds, or equivalently
@@
where $\epsilon_{N}=O(\frac{1}{\sqrt{N}})$. To establish this, we introduce two lemmas. Lemma 4.2 characterizes the coercivity condition. To facilitate a clearer proof, we first present Lemma 4.1.
@@
For any $i=1,2,\ldots,N$ and $v\in\mathcal{A}_{i}$, we construct the process
@@
Then, the following result
@@
holds.
We now proceed to formulate the coercivity condition rigorously, as stated in Lemma 4.2.
@@
Assume that Assumptions $\boldsymbol{(A_{2})}$ and $\boldsymbol{(A_{4})}$ hold. Then there exists constants $\epsilon>0,\ \epsilon^{\prime}>0$, and $C\in\mathbb{R}$, all independent of $N$ and $i$, such that
@@
Having established the necessary preparations, we now prove Theorem 4.5.
@@
Let $(\check{X}_{t}^{i},\ \check{P}_{t},\ \check{Q}_{t})$ denote the processes induced by the optimal control in (4.15). Under this notation, the difference in the social cost can be written as
@@
Applying Itô’s formula, we convert the terminal contribution into the integral form,
@@
Substituting this into (4.17) gives
@@
We now provide a more detailed explanation of how we derive $I_{1}$ and what $I_{2}$ characterizes. To derive $I_{1}$, we approximate $\sum_{i=1}^{N}\check{v}_{t}^{i}\Delta P_{t}$ by $\sum_{i=1}^{N}\overline{q}(t)\Delta P_{t}$ and approximate $\sum_{i=1}^{N}\check{P}_{t}\Delta v_{t}^{i}$ by $\sum_{i=1}^{N}\overline{p}(t)\Delta v_{t}^{i}$. We then use (4.18) to rewrite the terminal terms, and apply the optimal control formula (4.15) to eliminate $\check{v}_{t}^{i}$. This yields $I_{1}$, which represents the leading-order change in $J_{soc}$ caused by deviating from the proposed strategies. The second part, $I_{2}$, captures the approximation error due to replacing $(\check{v}_{t}^{i},\check{P}_{t})$ with $(\bar{q},\bar{p})$.
Following exactly the same argument as in the proof of Lemma 3.1, we obtain the existence of an $O(1)$ constant $K^{\prime\prime}$ such that
@@
$|J_{i}(\check{v}^{i},\check{\boldsymbol{v}}^{-i})|\leq K^{\prime\prime},\quad i=1,2,\ldots,N,$
which further implies
@@
By the coercivity condition in Lemma 4.2, there exists $O(1)$ constant $M_{1}>0$ such that, if $\mathbb{E}\left[\int_{0}^{T}Q_{t}^{2}dt\right]>M_{1}$, we have
@@
This inequality directly gives us the desired result. Thus, it suffices to establish the desired inequality (4.16) under the regime $\mathbb{E}\left[\int_{0}^{T}Q_{t}^{2}dt\right]\leq M_{1}$. Indeed, we will show that
@@
We first prove the claim for the term $I_{1}$. The next step relies on the identity
@@
We sketch the argument. Applying Itô’s formula and using the evolution of l{l} in (4.10), we obtain
@@
$=\mathbb{E}\left[\int_{0}^{T}\left(\bar{q}(t)\Delta u_{t}^{i}dt+\alpha{l}(t)\Delta v_{t}^{i}\right)dt\right],$ (4.21)
with $\Delta{P_{t}}$ and $\Delta u_{t}^{i}$ satisfying,
@@
Since $\Delta Q_{t}$ is the average of $\Delta v_{t}^{i}$, it follows from these dynamics that
$\frac{1}{N}\sum_{i=1}^{N}\Delta u_{t}^{i}=-\Delta P_{t},$ (4.24)
and then
@@
where the first and the second step are derived from (4.24) and (4) respectively. This leads directly to (4.20). Substituting (4.20) into the expression for $I_{1}$ and applying Itô’s formula again, we obtain
@@
Applying Itô’s formula once more, we obtain
@@
Let
$\overline{\Delta X_{t}}:=\frac{1}{N}\sum_{i=1}^{N}\Delta X_{t}^{i}$
be the averaged deviation process. Note that
@@
By Lemma 4.1, we conclude that
@@
In order to use this inequality, we invoke Assumption $\boldsymbol{(A_{4})}$, and apply fundamental inequality to obtain
@@
Subsequently, the convexity of $\frac{\eta}{2}x^{2}$ leads to
@@
where Jensen’s inequality is used to replace $\sum_{i=1}^{N}(\Delta X_{t}^{i})^{2}$ by $N\overline{\Delta X_{t}}^{2}$, and $\sum_{i=1}^{N}(\Delta v_{t}^{i})^{2}$ by $N\Delta Q_{t}^{2}$. Since the right hand side of (4.25) is nonnegative, we conclude
@@
The proof for $I_{1}$ is completed, and it is straightforward to prove the estimate for $I_{2}$. Similar as in Theorem 3.6’s proof, we have
@@
Using the Cauchy–Schwarz and triangle inequalities, together with the bound $\|Q\|^{2}\leq M_{1}$, we obtain
@@
Moreover, solving (4.22) and applying Cauchy–Schwarz again yields
$\Delta P_{t}=-\alpha\int_{0}^{t}e^{-\alpha(t-s)}\Delta Q_{s}ds,$
and
@@
Therefore, combining (4.26) and (4.27), we conclude that